Find the periodic payment R required to accumulate a sum of S dollars over t years with interest earned at the rate of r%/year compounded m times a year. (Round your answer to the nearest cent.) S = 300,000, r = 2.7, t = 10, m = 6

My question relates to finite math about periodic payment and it's a question from webassign homework
Answer #1

S=300,000, r=2.7, t=10, m=6

compond interest formula

M = S(1+i)n

M = 300,000(1+0.027)^(60) = 459460.681


answered by: gavin
Answer #2

SOLUTION :


Following formula will be applicable for periodic payment $R to accumulate $S in t years at interest rate of r% per year compounded m times a year ;


S = R ( (1 + r / (100m) )^(mt) - 1 ) / ( r / (100m) )   (ANSWER).


For S = $300000 ; r = 2.7% ; t = 10 years ; m = 6 , we have :


=> 300000 = R ( (1 + 2.7 / 600)^(60) - 1 ) / (2.7 / 600)

=> 300000 = R * 68.704726

=> R = 300000/68.704726

=> R = 4366.51 ($) (ANSWER).

answered by: Tulsiram Garg